OpenStudy (anonymous):

Solve 8x^2 = 128 by factoring. Select one: a. {-4, 4} b. {-8, 8} Incorrect c. {-8, 8} Incorrect d. 8

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

Its incorrect ?

OpenStudy (anonymous):

its correct.

OpenStudy (anonymous):

It has the solutions $$\pm 8$$

OpenStudy (anonymous):

I already tried it and it was wrong!!

OpenStudy (anonymous):

Well its not.

OpenStudy (anonymous):

B and C are the same thing .

OpenStudy (anonymous):

I'm a try D.

OpenStudy (anonymous):

Since A doesn't work

OpenStudy (zzr0ck3r):

8x^2 = 128 x^2 = 16 x = +- 4

OpenStudy (zzr0ck3r):

its a

OpenStudy (anonymous):

to many powers of 2

OpenStudy (anonymous):

lol

OpenStudy (zzr0ck3r):

:)

OpenStudy (anonymous):

so its not d

OpenStudy (zzr0ck3r):

its a 8x^2 = 128 divide each side by 8 x^2 = 16 what are two numbers where if you square them you get 16?

OpenStudy (zzr0ck3r):

-4 and 4

OpenStudy (anonymous):

yeah ur right thanks :) @zzr0ck3r